Problem 2-Buying a Car You see a car that you absolutely must have. The price tag says $18,450, but you want to keep your payment low over a 4-year period. Since you want to make payments no larger than $250 at the end of every month, you must find a bank to finance a loan for you. The bank you find charges interest at a rate of 6.99% compounded monthly. Use this information to find the following: The amount you need to borrow from the bank. . The down payment you need to make. .
Answer #1

Solution Rate per month = 6.99%/12 = 0.5825%

Total number of payments in 4 years = 12*4 = 48

The maximum amount of installment = $250

So present value of the 48 installments of $250 =

= 250* P/A(0.5825%,48) = 250*41.7683 = $10442.073

Therefore the amount need to borrow = $10442

The down payment we need to make = $18450 - $10442 = $8008